An iron cube has the volume of 4.76 cm^3. If iron has a density of 7.680 x 10^3 kg/m^3, how many moles are in the cube?


1
Expert's answer
2021-02-23T04:35:36-0500

"ρ = 7.680 \times 10^3 \; kg/m^3 \\

= 7.680 \times 10^6 \;g/m^3 \\

= 7.680 \;g/cm^3"

Proportion:

7.680 g – 1 cm3

x – 4.76 cm3

"x = \frac{7.680 \times 4.76}{1} = 36.55 \;g"

M(Fe) = 55.84 g/mol

"n = \frac{m}{M} \\

= \frac{36.55}{55.84} = 0.65 \;mol"

Answer: 0.65 mol